The ratio of the length of the hypotenuse to the length of the shorter leg in a right triangle is 8:5. The hypotenuse measures 144 meters. Find the length of the longer leg.
Answer by richwmiller(17219)

8/5=144/s
8s=5*144
s=5*144/8=90
a^2+90^2=144^2
a=112.41